# DUKAE: DUal-level Knowledge Accumulation and Ensemble for Pre-Trained Model-Based Continual Learning

### v1

@@ Abstract
Figure 1: Different PTMCL methods compared with our method. (a) PTMCL methods with misalignment issue. Classification heads are learned in different feature subspaces which are defined by task-specific PEFT modules and then kept fixed. (b) PTMCL methods merely fine-tune feature network with initial task data. Classification heads are learned in same feature space but lack the accumulation of feature knowledge for future tasks. (c) Our method leverages both feature-level and decision-level knowledge accumulation. Each task-specific feature subspace has a corresponding aligned classifier and memory is consolidated through the ensemble of subspace classification results.
1 Introduction
@@
Continual learning seeks to incrementally acquire new knowledge while retaining previously learned information from a continuous data stream. Traditionally, continual learning involves model starting from a randomly initialized parameter space and progressively accumulating new knowledge. In recent years, with the widespread application of pre-trained model (PTM) in natural language processing (NLP) [6, 2] and computer vision [13, 25], there has been an increasing amount of research in the field of PTM-based continual learning (PTMCL). Benefiting from the powerful foundational knowledge of PTM, PTMCL methods are akin to standing on the shoulders of giants, where knowledge is accumulated on top of an advanced starting point. These approaches significantly outperform traditional continual learning methods, which start from scratch and accumulate knowledge incrementally.
@@
Current PTMCL approaches can generally be categorized into two types. The first type [33] treats the PTM’s parameters as a new starting point, updating whole PTM parameters continuously without any additional parameters. The second type [32, 31, 27, 8, 30], which is more commonly used, keeps the PTM’s parameters fixed and acquires new knowledge through Parameter-Efficient Fine-Tuning (PEFT) [7]. These PEFT-based methods fine-tune additional parameters to learn task-specific knowledge, preserving the strong representation capabilities of the PTM. Typically, they fine-tune a small portion of parameters for each task, caching these additional parameters to retain knowledge. During inference, memory for previous tasks is maintained either by selecting the most relevant PEFT modules corresponding to the inference sample [32, 31, 30] or by integrating across all task-specific PEFT modules [27].
@@
Despite their advantages, these methods face significant challenges. A major issue is that each task-specific classification head is learned within different feature spaces and remains fixed once learned [32, 31, 27, 8], leading misalignment of classification heads problem (see figure (a) in Fig. 1). The misalignment of classification heads across tasks results in an inconsistency for inter-task comparison, leading to misclassification and, consequently, increased forgetting. Some methods [36, 24] avoid the misalignment issue by limiting feature network learning to the initial task only. (see figure (b) in Fig. 1). However, these methods rely solely on the first task’s data to train the feature network, without leveraging subsequent data to enhance representation capabilities. Consequently, their performance is limited by the feature discrimination ability.
@@
Figure 2: Comparison between our naive base method and LAE method (which suffers from misalignment problem) across four datasets under a 10 tasks continual learning setting. The y-axis represents average accuracy after learning the last task. The comparison is performed under identical network architecture and parameters configurations, illustrating that addressing the misalignment issue effectively mitigates forgetting.
2 Related Work
@@
Continual Learning (CL), also known as incremental learning or lifelong learning, is a research area focused on enabling models to learn from a continuous data stream without catastrophic forgetting. It can be typically categorized into three scenarios: Class-Incremental Learning (CIL), Domain-Incremental Learning (DIL), and Task-Incremental Learning (TIL) [28, 5, 19], in which CIL is the most challenging and widely studied. Recent continual learning methods can be broadly classified into three methodological paradigms. Replay-based methods mitigate forgetting by storing and replaying past experiences during learning new tasks [26, 3]. Regularization-based approaches incorporate additional terms in the loss function to protect important parameters of previous tasks, thus maintaining old knowledge [21, 14]. Architectural strategies involve dynamically expanding the model or isolating parameters specific to each task to manage new information effectively while retaining prior knowledge [23, 1].
@@
Parameter-Efficient Fine-Tuning (PEFT) has emerged as a significant advancement in adapting PTMs to downstream tasks with minimal additional parameters. Adapter-Tuning [10] initially introduces this concept by inserting lightweight, learnable modules into pre-trained transformers. Prompt-Tuning [18] and Prefix-Tuning [20] introduce the idea of modifying input prompts or hidden tokens, achieving notable success in NLP tasks. In the realm of vision transformers, VPT [12] and AdapterFormer [4] extend these ideas to visual tasks. LoRA [11] proposes learning low-rank matrices updates to efficiently fine-tune large models, while SSF [22] focuses on scaling and shifting operations within the model for better adaptation. NOAH [35] leverages a neural architecture search algorithm to design optimal prompt modules for large vision models.
@@
Pre-trained model-based continual learning (PTMCL) is attracting growing attention as it enables rapid learning of new knowledge by leveraging the robust foundational knowledge provided by PTM. L2P [32] is the first to introduce PTM into continual learning. DualPrompt [31] uses complementary prompts for task-invariant and task-specific instructions to manage sequential learning. S-Prompts [30] applies independent prompting across domains with cross-entropy loss for training and K-NN for domain identification. CODA-Prompt [27] employs a weighting mechanism to generate prompts, enhancing end-to-end task sequence learning without rehearsal. ADAM [36] aggregates embeddings from pre-trained and adapted models for classifier construction. EASE [37] utilizes expandable feature subspaces with lightweight adapters, ensuring efficient model updating without conflict. SLCA [33] integrates slow learning and classifier alignment, improving continual learning by progressively reducing learning rates. RanPAC [24] prevents forgetting with training-free random projectors and class-prototype accumulation.
@@
Figure 3: Discriminative capability of different feature networks across four datasets. We split each dataset into 10 tasks and train classifiers for all tasks using fixed feature network $f(\cdot;\theta,\mathcal{P}_{t})$ learned with dataset of different tasks $t$. The $10\times 10$ heatmap represents task-wise classification accuracy, while the $1\times 10$ heatmap indicates the average accuracy across tasks for each feature network. The results reveal distinct task-specific discriminative patterns and significant variations in overall feature discrimination capabilities for each feature network.
3 Preliminary
Figure 4: Illustration of our method, exemplified through the learning process on Task 5. (a) Feature-level knowledge accumulation process. The auxiliary classification head $\bar{\varphi}$ and the SSL branch are employed to facilitate the learning of the PEFT module $\mathcal{P}_{5}$, which is then cached for cumulative learning. (b) Decision-level knowledge accumulation and ensemble process. Using current task data, feature representations of $D_{5}$ are extracted in each feature subspace. These are then combined with Gaussian distributions $G_{i,j}$ from prior tasks $j$ to train classifiers $\varphi_{i}$ specific to each feature subspace $i$. Finally, adaptive expertise ensemble is applied to the ensemble of final output.
3.1 Formulation of PTMCL
@@
Currently, most PTMCL methods primarily rely on PEFT techniques for new tasks adaptation by fine-tuning a small set of additional parameters while keeping the PTM fixed. These techniques enables the model to efficiently learn new task-specific knowledge without compromising the original PTM’s core representation capabilities. We denote the PTMCL model as $g(f(\cdot;\theta,\mathcal{P}),\varphi)$, where $f(\cdot;\theta,\mathcal{P})$ represents the feature network of the model, with $\theta$ being the parameters of the PTM and $\mathcal{P}$ the additional parameters of the PEFT module (e.g., Adapter [10], LoRA [11]). The classifier for the downstream tasks is denoted as $g(\cdot;\varphi)$, where $\varphi$ refers to the parameters of classifier. The PTM module parameters $\theta$ remain fixed during the continual learning, while the PEFT module parameters $\mathcal{P}$ and the classifier parameters $\varphi$ are incrementally fine-tuned as new tasks arrive. The mathematical formulation of PTMCL can be generalized as optimizing the following objective function:
$\min_{\mathcal{P},\varphi}\sum_{(x,y)\in D_{1:T}}\mathcal{L}\left(g\left(f\left(x;\theta,\mathcal{P}\right);\varphi\right),y\right)+\mathcal{L}_{\text{reg}}(\mathcal{P}),$ (1)
@@
3.2 Naive Base Method
Previous PTMCL methods primarily employ PEFT to accumulate task-specific feature-level knowledge. However, many methods [32, 31, 27, 8] often neglect the importance of aligning classification heads across tasks. Typically, once a task-specific classification head is trained, it remains fixed, leading to different tasks’ classification heads learned in different feature spaces. This misalignment of classification heads creates inconsistent decision boundaries across tasks, making it difficult to effectively compare classification results, consequently increasing forgetting.
@@ 3.2 Naive Base Method
$\begin{split}\min_{\varphi_{1}}&\left[\sum_{(x,y)\in D_{t}}\mathcal{L}\left(g\left(f(x;\theta,\mathcal{P}_{1});\varphi_{1}\right),y\right)\right.\\ &\left.+\sum_{c=1}^{|C_{1:t-1}|}\sum_{(x_{c},y_{c})\sim\mathcal{G}_{1,c}}\mathcal{L}\left(g(x_{c};\varphi_{1}),y_{c}\right)\right].\end{split}$ (2)
@@
To empirically verify the relationship between misalignment and catastrophic forgetting, we conducted a controlled comparative study between our naive base method and the LAE [8] approach (which suffers from misalignment) while maintaining identical network architecture and parameters configurations. The experimental results reveal that mitigating misalignment effectively alleviates forgetting and enhances overall performance (see Fig. 2).
4 DUal-level Knowledge Accumulation and Ensemble
@@
4.1 Feature-level Knowledge Accumulation
To achieve the feature-level knowledge accumulation, we use an expansion strategy same as [37] to store new task-specific PEFT modules. Typically, for each subsequent task $t$, a new PEFT module $\mathcal{P}_{t}$ is trained using the current task’s data $D_{t}$, and we cache the task-specific PEFT modules for all seen tasks $\{\mathcal{P}_{1},\ldots,\mathcal{P}_{t}\}$. For each task $t$, we train the associated PEFT module $\mathcal{P}_{t}$ using a cross-entropy loss function:
$\min_{\mathcal{P}_{t},\bar{\varphi}}\sum_{(x,y)\in D_{t}}\mathcal{L}_{CE}\left(g(f(x;\theta,\mathcal{P}_{t});\bar{\varphi}),y\right),$ (3)
@@ 4.1 Feature-level Knowledge Accumulation
To further boost feature discrimination ability, we integrate an auxiliary self-supervised learning (SSL) branch [17, 38]. This branch simply classify four new classes by rotating input data by 0, 90, 180, and 270 degrees. The SSL loss is defined as:
$\mathcal{L}_{SSL}=\mathcal{L}_{CE}\left(g(f(Rot(x);\theta,\mathcal{P}_{t});\bar{\varphi}_{SSL}),Rot(y)\right),$ (4)
@@
$\mathcal{L}_{\mathcal{P}_{t}}=\mathcal{L}_{CE}\left(g(f(x;\theta,\mathcal{P}_{t});\bar{\varphi}),y\right)+\alpha\cdot\mathcal{L}_{SSL},$ (5)
@@
4.2 Decision-level Knowledge Accumulation

### v2

@@ Abstract.
Figure 1. Different PTMCL methods compared with our method. (a) PTMCL methods with misalignment issue. Classification heads are learned in different feature subspaces which are defined by task-specific PEFT modules and then kept fixed. (b) PTMCL methods merely fine-tune feature network with initial task data. Classification heads are learned in same feature space but lack the accumulation of feature knowledge for future tasks. (c) Our method leverages both feature-level and decision-level knowledge accumulation. Each task-specific feature subspace has a corresponding aligned classifier and memory is consolidated through the ensemble of subspace classification results.
1. Introduction
@@
Continual learning seeks to incrementally acquire new knowledge while retaining previously learned information from a continuous data stream. Traditionally, continual learning involves model starting from a randomly initialized parameter space and progressively accumulating new knowledge. In recent years, with the widespread application of pre-trained model (PTM) in natural language processing (NLP) (Devlin et al., 2018; Brown et al., 2020) and computer vision (Kirillov et al., 2023; Radford et al., 2021), there has been an increasing amount of research in the field of PTM-based continual learning (PTMCL). Benefiting from the powerful foundational knowledge of PTM, PTMCL methods are akin to standing on the shoulders of giants, where knowledge is accumulated on top of an advanced starting point. These approaches significantly outperform traditional continual learning methods, which start from scratch and accumulate knowledge incrementally.
@@
Current PTMCL approaches can generally be categorized into two types. The first type (Zhang et al., 2023) treats the PTM’s parameters as a new starting point, updating whole PTM parameters continuously without any additional parameters. The second type (Wang et al., 2022c, b; Smith et al., 2023; Gao et al., 2023; Wang et al., 2022a), which is more commonly used, keeps the PTM’s parameters fixed and acquires new knowledge through Parameter-Efficient Fine-Tuning (PEFT) (Ding et al., 2023). These PEFT-based methods fine-tune additional parameters to learn task-specific knowledge, preserving the strong representation capabilities of the PTM. Typically, they fine-tune a small portion of parameters for each task, caching these additional parameters to retain knowledge. During inference, memory for previous tasks is maintained either by selecting the most relevant PEFT modules corresponding to the inference sample (Wang et al., 2022c, b, a) or by integrating across all task-specific PEFT modules (Smith et al., 2023).
@@
Despite their advantages, these methods face significant challenges. A major issue is that each task-specific classification head is learned within different feature spaces and remains fixed once learned (Wang et al., 2022c, b; Smith et al., 2023; Gao et al., 2023), leading misalignment of classification heads problem (see figure (a) in Fig. 1). The misalignment of classification heads across tasks results in an inconsistency for inter-task comparison, leading to misclassification and, consequently, increased forgetting. Some methods (Zhou et al., 2024a; McDonnell et al., 2024) avoid the misalignment issue by limiting feature network learning to the initial task only. (see figure (b) in Fig. 1). However, these methods rely solely on the first task’s data to train the feature network, without leveraging subsequent data to enhance representation capabilities. Consequently, their performance is limited by the feature discrimination ability.
@@
Figure 2. Comparison between our naive base method and LAE method (which suffers from misalignment problem) across four datasets under a 10 tasks continual learning setting. The y-axis represents average accuracy after learning the last task. The comparison is performed under identical network architecture and parameters configurations, illustrating that addressing the misalignment issue effectively mitigates forgetting.
2. Related Work
@@
Continual Learning (CL), also known as incremental learning or lifelong learning, is a research area focused on enabling models to learn from a continuous data stream without catastrophic forgetting. It can be typically categorized into three scenarios: Class-Incremental Learning (CIL), Domain-Incremental Learning (DIL), and Task-Incremental Learning (TIL) (Van de Ven and Tolias, 2019; De Lange et al., 2021; Li et al., 2024), in which CIL is the most challenging and widely studied. Recent continual learning methods can be broadly classified into three methodological paradigms. Replay-based methods mitigate forgetting by storing and replaying past experiences during learning new tasks (Rebuffi et al., 2017; Castro et al., 2018). Regularization-based approaches incorporate additional terms in the loss function to protect important parameters of previous tasks, thus maintaining old knowledge (Li and Hoiem, 2017; Kirkpatrick et al., 2017). Architectural strategies involve dynamically expanding the model or isolating parameters specific to each task to manage new information effectively while retaining prior knowledge (Mallya and Lazebnik, 2018; Aljundi et al., 2017).
@@
Parameter-Efficient Fine-Tuning (PEFT) has emerged as a significant advancement in adapting PTMs to downstream tasks with minimal additional parameters. Adapter-Tuning (Houlsby et al., 2019) initially introduces this concept by inserting lightweight, learnable modules into pre-trained transformers. Prompt-Tuning (Lester et al., 2021) and Prefix-Tuning (Li and Liang, 2021) introduce the idea of modifying input prompts or hidden tokens, achieving notable success in NLP tasks. In the realm of vision transformers, VPT (Jia et al., 2022) and AdapterFormer (Chen et al., 2022) extend these ideas to visual tasks. LoRA (Hu et al., 2021) proposes learning low-rank matrices updates to efficiently fine-tune large models, while SSF (Lian et al., 2022) focuses on scaling and shifting operations within the model for better adaptation. NOAH (Zhang et al., 2024b) leverages a neural architecture search algorithm to design optimal prompt modules for large vision models.
@@
Pre-trained model-based continual learning (PTMCL) is attracting growing attention as it enables rapid learning of new knowledge by leveraging the robust foundational knowledge provided by PTM. L2P (Wang et al., 2022c) is the first to introduce PTM into continual learning. DualPrompt (Wang et al., 2022b) uses complementary prompts for task-invariant and task-specific instructions to manage sequential learning. S-Prompts (Wang et al., 2022a) applies independent prompting across domains with cross-entropy loss for training and K-NN for domain identification. CODA-Prompt (Smith et al., 2023) employs a weighting mechanism to generate prompts, enhancing end-to-end task sequence learning without rehearsal. ADAM (Zhou et al., 2024a) aggregates embeddings from pre-trained and adapted models for classifier construction. EASE (Zhou et al., 2024b) utilizes expandable feature subspaces with lightweight adapters, ensuring efficient model updating without conflict. SLCA (Zhang et al., 2023) integrates slow learning and classifier alignment, improving continual learning by progressively reducing learning rates. RanPAC (McDonnell et al., 2024) prevents forgetting with training-free random projectors and class-prototype accumulation.
@@
Figure 3. Discriminative capability of different feature networks across four datasets. We split each dataset into 10 tasks and train classifiers for all tasks using fixed feature network $f(\cdot;\theta,\mathcal{P}_{t})$ learned with dataset of different tasks $t$. The $10\times 10$ heatmap represents task-wise classification accuracy, while the $1\times 10$ heatmap indicates the average accuracy across tasks for each feature network. The results reveal distinct task-specific discriminative patterns and significant variations in overall feature discrimination capabilities for each feature network.
3. Preliminary
3.1. Formulation of PTMCL
@@
Currently, most PTMCL methods primarily rely on PEFT techniques for new tasks adaptation by fine-tuning a small set of additional parameters while keeping the PTM fixed. These techniques enables the model to efficiently learn new task-specific knowledge without compromising the original PTM’s core representation capabilities. We denote the PTMCL model as $g(f(\cdot;\theta,\mathcal{P}),\varphi)$, where $f(\cdot;\theta,\mathcal{P})$ represents the feature network of the model, with $\theta$ being the parameters of the PTM and $\mathcal{P}$ the additional parameters of the PEFT module (e.g., Adapter (Houlsby et al., 2019), LoRA (Hu et al., 2021)). The classifier for the downstream tasks is denoted as $g(\cdot;\varphi)$, where $\varphi$ refers to the parameters of classifier. The PTM module parameters $\theta$ remain fixed during the continual learning, while the PEFT module parameters $\mathcal{P}$ and the classifier parameters $\varphi$ are incrementally fine-tuned as new tasks arrive. The mathematical formulation of PTMCL can be generalized as optimizing the following objective function:
(1) $\min_{\mathcal{P},\varphi}\sum_{(x,y)\in D_{1:T}}\mathcal{L}\left(g\left(f\left(x;\theta,\mathcal{P}\right);\varphi\right),y\right)+\mathcal{L}_{\text{reg}}(\mathcal{P}),$
@@
Figure 4. Illustration of our method, exemplified through the learning process on Task 5. (a) Feature-level knowledge accumulation process. The auxiliary classification head $\bar{\varphi}$ and the SSL branch are employed to facilitate the learning of the PEFT module $\mathcal{P}_{5}$, which is then cached for cumulative learning. (b) Decision-level knowledge accumulation and ensemble process. Using current task data, feature representations of $D_{5}$ are extracted in each feature subspace. These are then combined with Gaussian distributions $G_{i,j}$ from prior tasks $j$ to train classifiers $\varphi_{i}$ specific to each feature subspace $i$. Finally, adaptive expertise ensemble is applied to the ensemble of final output.
3.2. Naive Base Method
Previous PTMCL methods primarily employ PEFT to accumulate task-specific feature-level knowledge. However, many methods (Wang et al., 2022c, b; Smith et al., 2023; Gao et al., 2023) often neglect the importance of aligning classification heads across tasks. Typically, once a task-specific classification head is trained, it remains fixed, leading to different tasks’ classification heads learned in different feature spaces. This misalignment of classification heads creates inconsistent decision boundaries across tasks, making it difficult to effectively compare classification results, consequently increasing forgetting.
@@ 3.2. Naive Base Method
(2) $\begin{split}\min_{\varphi_{1}}&\left[\sum_{(x,y)\in D_{t}}\mathcal{L}\left(g\left(f(x;\theta,\mathcal{P}_{1});\varphi_{1}\right),y\right)\right.\\ &\left.+\sum_{c=1}^{|C_{1:t-1}|}\sum_{(x_{c},y_{c})\sim\mathcal{G}_{1,c}}\mathcal{L}\left(g(x_{c};\varphi_{1}),y_{c}\right)\right].\end{split}$
@@
To empirically verify the relationship between misalignment and catastrophic forgetting, we conducted a controlled comparative study between our naive base method and the LAE (Gao et al., 2023) approach (which suffers from misalignment) while maintaining identical network architecture and parameters configurations. The experimental results reveal that mitigating misalignment effectively alleviates forgetting and enhances overall performance (see Fig. 2).
4. DUal-level Knowledge Accumulation and Ensemble
Table 1. Result of LAA and IAA for baseline methods and our approach across four datasets. * means results from our re-implementation.
Method CIFAR-100 ImageNet-R CUB-200 Cars-196
LAA (%) IAA (%) LAA (%) IAA (%) LAA (%) IAA (%) LAA (%) IAA (%)
L2P (Wang et al., 2022c) 82.76±1.17 88.48±0.83 66.49±0.40 72.83±0.56 62.21±1.92 73.83±1.67 38.18±2.33 51.79±4.19
DualPrompt (Wang et al., 2022b) 85.56±0.33 90.33±0.33 68.50±0.52 72.59±0.24 66.00±0.57 77.92±0.50 40.14±2.36 56.74±1.78
CODA-Prompt (Smith et al., 2023) 86.56±0.77 90.61±0.36 75.25±0.56 81.26±0.76 72.63±0.76 80.54±0.54 44.89±0.61 58.91±0.37
LAE (Gao et al., 2023) 85.59±0.46 89.96±0.44 72.66±0.63 78.91±0.89 77.48±0.94 85.83±0.68 52.47±1.46 64.08±1.01
ADAM (Zhou et al., 2024a) 87.46±0.03 92.20±0.06 66.70±0.21 75.18±0.07 86.77±0.02 91.32±0.01 41.77±7.73 53.78±7.76
EASE (Zhou et al., 2024b) 87.73±0.17 92.29±0.19 75.89±0.28 81.67±0.20 86.62±0.08 91.20±0.03 37.58±0.25 51.00±0.14
SLCA (Zhang et al., 2023) 91.53±0.28 94.09±0.87 77.00±0.33 81.17±0.64 84.71±0.40 90.94±0.68 67.73±0.85 76.93±1.21
SLCA++ (Zhang et al., 2024a) 91.69±0.15 94.47±0.72 79.78±0.16 84.31±0.73 86.59±0.29 91.63±0.72 73.97±0.22 79.46±0.80
RanPAC (McDonnell et al., 2024) 92.10±0.17 95.13±0.06 77.44±0.34 83.06±0.30 89.09±0.23 92.86±0.03 74.69±0.92 82.24±0.41
@@ 4. DUal-level Knowledge Accumulation and Ensemble
4.1. Feature-level Knowledge Accumulation
To achieve the feature-level knowledge accumulation, we use an expansion strategy same as (Zhou et al., 2024b) to store new task-specific PEFT modules. Typically, for each subsequent task $t$, a new PEFT module $\mathcal{P}_{t}$ is trained using the current task’s data $D_{t}$, and we cache the task-specific PEFT modules for all seen tasks $\{\mathcal{P}_{1},\ldots,\mathcal{P}_{t}\}$. For each task $t$, we train the associated PEFT module $\mathcal{P}_{t}$ using a cross-entropy loss function:
(3) $\min_{\mathcal{P}_{t},\bar{\varphi}}\sum_{(x,y)\in D_{t}}\mathcal{L}_{CE}\left(g(f(x;\theta,\mathcal{P}_{t});\bar{\varphi}),y\right),$
@@ 4.1. Feature-level Knowledge Accumulation
To further boost feature discrimination ability, we integrate an auxiliary self-supervised learning (SSL) branch (Lee et al., 2020; Zhu et al., 2021). This branch simply classify four new classes by rotating input data by 0, 90, 180, and 270 degrees. The SSL loss is defined as:
(4) $\mathcal{L}_{SSL}=\mathcal{L}_{CE}\left(g(f(Rot(x);\theta,\mathcal{P}_{t});\bar{\varphi}_{SSL}),Rot(y)\right),$
@@
(5) $\mathcal{L}_{\mathcal{P}_{t}}=\mathcal{L}_{CE}\left(g(f(x;\theta,\mathcal{P}_{t});\bar{\varphi}),y\right)+\alpha\cdot\mathcal{L}_{SSL},$
@@
4.2. Decision-level Knowledge Accumulation